OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=7tan(4x−pi/7)

OpenStudy (anonymous):

First two non negative asymptotes: x = 9/56 pi radians and then x =23/56 pi radians. first negative asymptote is x = -5/56 pi radians